OpenStudy (anonymous):

1 2/13 divided by 1/13?? should be interger or simplified fraction

OpenStudy (anonymous):

yes you better convert the "mixed number" to an improper fraction first

OpenStudy (anonymous):

yes you better convert the "mixed number" to an improper fraction first

myininaya (myininaya):

\[\frac{13 \cdot 1 +2}{13} \cdot \frac{13}{1}=13 \cdot 1+2\]

myininaya (myininaya):

=15

OpenStudy (anonymous):

\[1\tfrac{2}{13}=\frac{15}{13}\] then use \[\frac{15}{13}\times \frac{13}{1}\] to find the answer

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so its 195/13??

myininaya (myininaya):

why would you multiply 15 and 13 when there is 13 on top and 13 on bottom you can go ahead in cancel the 13's like i did above

OpenStudy (anonymous):

so it would be 15/169

myininaya (myininaya):

\[\frac{15}{13} \cdot \frac{13}{1} =\frac{15}{\not{13}} \cdot \frac{\not{13}}{1}\]

OpenStudy (anonymous):

nevermind i got it 15 was the answer

myininaya (myininaya):

what do you have left on top?

OpenStudy (anonymous):

15

myininaya (myininaya):

right and you have 1 left on bottom so 15/1=15

OpenStudy (anonymous):

much appericated